OpenStudy (anonymous):

5/6c=7/24 what is c??

OpenStudy (anonymous):

Hello anyone gonna help??

OpenStudy (precal):

Multiply both sides by (6/5)

OpenStudy (anonymous):

If you cross multiply you get: 120 = 42c Then divide each side by 42. You get 120/42.

OpenStudy (precal):

c=(6/5)(7/24) c=7/35

OpenStudy (precal):

I mean (7/20)

OpenStudy (anonymous):

If you multiply each side by (6/5) you get 1/c = 42/120, or c = 120/42

OpenStudy (precal):

I am assuming c is in the numerator. Most people do not use equation tab to enter in the equations.

OpenStudy (anonymous):

You're right it totally depends on if it's (5/6)c or 5/(6c) so I guess it's up to '2cool4skool' to figure the rest out.